# https://www.usc.edu/

### v1

@@ 7 Evaluation Metrics
[…]
    where $n$ is the number of time steps, $p_{i}$ and $q_{i}$ are the positions at time step i in the estimated and ground truth trajectories, respectively, and $p_{i}-q_{i}||$ is the Euclidean distance between $p_{i}$ and $q_{i}$.
@@
[…]
    where n is the number of time steps, $\Delta p_{i}$ and $\Delta q_{i}$ are the relative pose changes between time steps $i$ and $i+\Delta$ in the estimated and ground truth trajectories, respectively, and $\Delta p_{i}-\Delta q_{i}||$ is the error metric used to measure the difference between the relative poses (e.g., Euclidean distance for position errors, angular distance for orientation errors).
@@
[…]

### v2

@@ 7 Evaluation Metrics
    $ATE=sqrt\left(\frac{1}{n}*\sum||p_{i}-q_{i}||^{2}\right)$ (92)
    where $n$ is the number of time steps, $p_{i}$ and $q_{i}$ are the positions at time step i in the estimated and ground truth trajectories, respectively, and $||p_{i}-q_{i}||$ is the Euclidean distance between $p_{i}$ and $q_{i}$.
@@
    $RPE=sqrt\left(\frac{1}{n-\Delta}*\sum(||\Delta p_{i}-\Delta q_{i}||^{2})\right)$ (93)
    where n is the number of time steps, $\Delta p_{i}$ and $\Delta q_{i}$ are the relative pose changes between time steps $i$ and $i+\Delta$ in the estimated and ground truth trajectories, respectively, and $||\Delta p_{i}-\Delta q_{i}||$ is the error metric used to measure the difference between the relative poses (e.g., Euclidean distance for position errors, angular distance for orientation errors).
@@
    $\textrm{AVI}=\frac{1}{n}\sum_{r\in\hat{i}_{sk}^{j}(r)}{\textrm{Mask}_{j,j+1}\cdot||\textrm{Warp}(\hat{i}_{sk}^{j})(r)-\hat{i}_{sk}^{j+1}(r)||_{2}}$ (94)